OpenStudy (anonymous):

a=-2 , b=-1 , c=4 , d=3 Solve) l a l - l b l = ?

OpenStudy (anonymous):

1

OpenStudy (lgbasallote):

the | | means absolute value which means take the positive so 2 - 1

OpenStudy (anonymous):

-2=2 -1=1 2-1=1